# How Much are Large Language Models Contaminated? A Comprehensive Survey and the LLMSanitize Library

### v3

@@ 3.4 Model-based Closed-Data Contamination Detection
Through the use case of fine-tuning LLMs on mathematical reasoning datasets like GSM8K (Cobbe et al., 2021), Tu et al. (2024) introduce a new paradigm in (closed-data) contamination detection. Their DICE approach consists in finding the layer of the fine-tuned LLM furthest away from its counterpart layer in the not fine-tuned same LLM. Then, a MLP binary classifier is inferred on the isolated layer’s weights to predict contamination. This method is particularly effective for in-distribution contamination detection, a setup defined by the authors as fine-tuning on a dataset from the same distribution as the evaluation dataset. In this setup, DICE reaches near-perfect AUC, outperforming all other methods, including notably Min-KProb.
@@ 4 Discussion
4.1 Best Practices
@@ 4.2 New Evaluation Benchmarks
Several datasets have been introduced aiming to provide contamination-free evaluation.
@@
Some new datasets discard the static paradigm of existing LLM evaluation datasets, and instead propose a dynamic benchmark to avoid contamination:
@@ 4.3 Future Challenges
We believe in the need for comprehensive ethical and legal frameworks (Chang, 2021) to govern the collection, usage, and management of data for LLM training. This includes policies and protocols for data privacy, consent, and use that help prevent the incorporation of contaminated data from unethical sources and the contamination of widely used pre-training data sources (e.g., CommonCrawl). Developing contamination detection techniques that do not compromise individual privacy will be crucial (Lin et al., 2010; Hayes and Ohrimenko, 2018).
5 The LLMSanitize Library
LLM ARC HellaSwag MMLU TruthulQA Winogrande GSM8K
Guided Prompting (ROUGE-L)
Llama-2-7B 0.064 0.489 0.146 0.001 0.001 0.968
Qwen-1.5-7B 0.000 0.023 0.700 0.664 0.202 1.000
Mistral-7B 0.000 0.000 0.010 0.042 0.999 0.444
Gemma-7B 0.978 0.000 0.580 0.122 0.820 0.000
GPT-3.5-turbo 0.000 0.184 0.970 0.075 0.000 1.000
Sharded Likelihood
Llama-2-7B 0.258 0.000 0.967 0.154 0.171 0.781
Qwen-1.5-7B 0.426 0.000 0.476 0.255 0.115 0.209
Mistral-7B 0.862 0.000 0.376 0.057 0.485 0.507
Gemma-7B 0.234 0.000 0.255 0.450 0.215 0.648
Table 1: P-value for two closed-data contamination detection methods using a statistical test: Guided Prompting (Golchin and Surdeanu, 2023a) (top) and Sharded Likelihood (Oren et al., 2023) (bottom). Shaded results are not statistically significant. We ran LLMSanitize on 100 test data points.
To facilitate progress in contamination detection in LLMs, we have built and are releasing an open-source Python library, named LLMSanitize. LLMSanitize supports most of the open-data and closed-data contamination detection described above and listed in Figure 1. It relies on the HuggingFace hub to access LLMs and the data, namely transformers (Wolf et al., 2020) and datasets (Lhoest et al., 2021) libraries. Efficient LLM inference for model-based use cases is handled with the vLLM library (Kwon et al., 2023). LLMSanitize is straightforward to install through PyPI. Users are expected to pass in the relevant arguments for each contamination detection method, as well as key parameters such as the LLM and dataset name and the number of data points to run the contamination detection on. We refer to Appendix A.
We now put to test LLMSanitize. We apply Guided Prompting (Golchin and Surdeanu, 2023b) and Sharded Likelihood (Oren et al., 2023) on the six datasets of HuggingFace OpenLLM Leaderboard [^1] https://huggingface.co/open-llm-leaderboard. We employ four widely used 7B-size LLMs: Llama-2 (Touvron et al., 2023), Qwen-1.5 (Bai et al., 2023), Mistral (Jiang et al., 2023) and Gemma (Gemma Team Google DeepMind, 2024). We also evaluate GPT-3.5-turbo on guided prompting - the blocked access to the logprobs prohibits us from applying the other method. Results are shown in Table 1. The HellaSwag benchmark (Zellers et al., 2019) shows strong signs of contamination, consistent across methods and LLMs.
@@ 6 Conclusion
As LLMs rapidly evolve and the training data become more extensive, LLMs’ performances are inevitably biased by contamination. Therefore, in this survey, we investigated and structured the current landscape of contamination detection in LLMs. We introduced a new Python library named LLMSanitize which is a universal tool for quickly detecting contamination in various models and datasets. The extensive investigation and library could provide a universal foundation for future models when considering contamination issues.
Limitations
In this paper, we reviewed all work related to the detection of contamination. Contamination itself, as a general problem with LLMs, is larger than detection. For instance, our study scope does not focus on the mitigation of contamination, which we only briefly touch in Section 4.1. Mitigation is also of critical importance ; but we made the choice in this study to narrow down and focus on detection, which we believe makes our contribution more focused.
We included all works mentioning the issue of detecting or assessing contamination with language models from the last four years. Many of these works, for instance language models reports (Radford et al., 2019), are not focusing on the issue of contamination detection: these papers simply discuss this issue in a few paragraphs. Our survey also includes papers entirely dedicated to contamination detection, such as the LLM Decontaminator (Yang et al., 2023b). The technical depth of the work related to contamination detection across such differing types of works is very heterogeneous. This is not reflected in our survey as we chose to include all papers on the same level.
Appendix A LLMSanitize Usage
In this section we showcase how to use our Python library LLMSanitize. The library is recommended to be used with Python 3.9 and CUDA 11.8. The first step is to create the environment and install the library.
The following code should be run in a terminal:
1conda create --name llmsanitize python=3.9
2conda activate llmsanitize
3pip install llmsanitize==0.0.8
We first review how to use LLMSanitize for an open-data contamination detection setup. In this case, we compare two datasets and no LLM is involved. Let us take a simple setup: we compare HellaSwag as evaluation dataset against ARC as training dataset, sub-sampling 100 evaluation data points, and using GPT-2’s 8-grams collision as contamination detection method.
The following Python code could be run:
1import argparse
2from llmsanitize import OpenDataContaminationChecker
3
4parser = argparse.ArgumentParser()
5parser.add_argument("--train_data_name", type=str, default="allenai/ai2_arc")
6parser.add_argument("--eval_data_name", type=str, default="Rowan/hellaswag")
7parser.add_argument("--eval_set_key", type=str, default="validation")
8parser.add_argument("--text_key", type=str, default="ctx")
9parser.add_argument("--n_eval_data_points", type=int, default=100)
10parser.add_argument("--method", type=str, default="gpt-2")
11args = parser.parse_args()
12
13contamination_checker = OpenDataContaminationChecker(args)
14contamination_checker.run_contamination(args.method)
From there, the library will run the contamination detection and log results to a dedicated file. We refer the reader to all argument fields defined in the main.py file ; which among other properties, provide an option to stream over the (large) training dataset in order to limit memory usage for the use case of open-data contamination.
Next, we review how to conduct a closed-data contamination detection use case, involving an LLM and an evaluation dataset. For this setup, we apply the Guided Prompting method on Llama-2 on the ARC dataset. For this method (and all closed-data contamination detection methods except Sharded Likelihood), we make use of vLLM for efficient inference.
This requires us to first open a vLLM instance. Run the following shell script, which assumes access to 2 GPUs:
1export LD_LIBRARY_PATH=/usr/local/cuda/lib64:\$LD_LIBRARY_PATH
2export RAY_memory_monitor_refresh_ms=0;
3export CUDA_VISIBLE_DEVICES=0,1;
4server_type=vllm.entrypoints.openai.api_server
5
6python -m \$server_type \
7 --model <path_to_local_llama2> \
8 --tokenizer <path_to_local_llama2> \
9 --gpu-memory-utilization=0.9 \
10 --max-num-seqs=200 \
11 --disable-log-requests \
12 --host 127.0.0.1 \
13 --port <port_number> \
14 --tensor-parallel-size 2 \
15 --download-dir /export/home/cache
Next, a minimal Python code to run Guided Prompting detection could be:
1import argparse
2from llmsanitize import ClosedDataContaminationChecker
3
4parser = argparse.ArgumentParser()
5parser.add_argument("--eval_data_name", type=str, default="allenai/ai2_arc")
6parser.add_argument("--eval_set_key", type=str, default="test")
7parser.add_argument("--text_key", type=str, default="question")
8parser.add_argument("--label_key", type=str, default="anskwerKey")
9parser.add_argument("--n_eval_data_points", type=int, default=100)
10parser.add_argument("--num_proc", type=int, default=100)
11parser.add_argument("--method", type=str, default="guided-prompting")
12parser.add_argument("--local_port", type=int, default=<port_number>)
13parser.add_argument("--model_name", type=str, default=<path_to_local_llama2>)
14parser.add_argument("--guided_prompting_task_type", type=str, default="CLS")
15parser.add_argument("--user_local_model", type=bool, default=True)
16args = parser.parse_args()
17
18contamination_checker = ClosedDataContaminationChecker(args)
19contamination_checker.run_contamination(args.method)
Note that the model path and port number passed as arguments in the Python script should match the ones used to launch the vLLM instance.

### v4

@@ 3.4.2 Probing Hidden Layers
Through the use case of fine-tuning LLMs on mathematical reasoning datasets like GSM8K (Cobbe et al., 2021), Tu et al. (2024) introduce a new paradigm in (closed-data) contamination detection. Their DICE approach consists in finding the layer of the fine-tuned LLM furthest away from its counterpart layer in the not fine-tuned same LLM. Then, a MLP binary classifier is inferred on the isolated layer’s weights to predict contamination. This method is particularly effective for in-distribution contamination detection, a setup defined by the authors as fine-tuning on a dataset from the same distribution as the evaluation dataset. In this setup, DICE reaches near-perfect ROC-AUC, outperforming all other methods, including notably Min-KProb.
Closed-Data Method Type Black-box Gray-box White-box
Performance Analysis CAP Zhao et al. (2024)
Aiyappa et al. (2023)
Roberts et al. (2023)
Li & Flanigan (2023)
Cao et al. (2024)
Membership Inference Attacks RECALL (Xie et al., 2024)
MIA-Tuner (Fu et al., 2024a)
MaxNorm(Meeus et al., 2024a)
PREMIA (Feng et al., 2024)
SPV-MIA (Fu et al., 2023)
CAMIA (Chang et al., 2024)
S2-MIA (Li et al., 2024)
EM-MIA (Kim et al., 2024)
Model Memorization Gemma Team Google DeepMind (2024)
Magar & Schwartz (2022)
Nasr et al. (2023)
According To prompting (Weller et al., 2023)
DCQ (Samuel et al., 2024)
Local order quiz (Samuel et al., 2024)
TS-Guessing (Deng et al., 2023)
Chang et al. (2023)
Ranaldi et al. (2024)
Model Confidence Min-KProb (Shi et al., 2024)
CDD and TED (Dong et al., 2024)
DC-PDD (Zhang et al., 2024b)
Min-K%++ (Zhang et al., 2024a)
SURP (Zhang & Wu, 2024)
Li (2023a)
Oren et al. (2023)
Xu et al. (2024b)
Liu et al. (2024)
DICE (Tu et al., 2024)
Table 2: Classification of closed-data contamination detection methods as per the level of access to the LLM required.
Liu et al. (2024) also make use of the intermediate layers of a fine-tuned LLM. Their method first gathers a dataset of non-members and members ; where the former are pulled from after the pre-training cutoff or synthetically generated from ChatGPT. Then, the LLM is fine-tuned to classify membership between these two groups. Activations of intermediate layers of the fine-tuned LLM on both members and non-members are used as input of a probe classifier assessing membership. At inference time, the whole system feeds a target text to the fine-tuned LLM, extract its activations and feeds them to the probe classifier to check membership. The best layer to extract activation from is determined on a validation set. Results show greater ROC-AUC on WikiMIA (Shi et al., 2024) and ArxivMIA (Liu et al., 2024), and indicate that this probing system works better with larger models.
Table 2 summarizes closed-data detection techniques organized by the minimum level of access to the LLM which is required. All methods based on performance analysis only require a black-box access, making them convenient to use. Membership-inference attacks are equally split between black-box and gray-box access requirements. Model memorization techniques mainly rely on prompting and thus only need black-box access. Lastly, for model confidence methods, which often analyze output distributions, a gray-box or even white-box model access is needed.
@@ 4 Discussion
4.1 Best Practices to Avoid Contamination
@@ 4.2 New Evaluation Benchmarks
Researchers have crafted new datasets specifically introduced with the aim to provide contamination-free LLM evaluation.
Open-Source Pre-training Data
When the pre-training data is publicly available, the task of defining members and non-members of the pre-training set becomes drastically simpler. Duan et al. (2024) introduce the MIMIR dataset, constructed from The Pile training set (Gao et al., 2020), on the which popular open-source models such the Pythia (Biderman et al., 2023) and GPT-Neo [^2] https://github.com/EleutherAI/gpt-neo series are pre-trained. MIMIR covers domain-specific subsets (e.g. GitHub, Wikipedia, Arxiv) and extracts members through several n-gram sizes, allowing researchers to work on multiple levels of membership granularity. MIMIR has become widely used in MIA research with LLMs. OLMoMIA (Kim et al., 2024) introduces a membership inference dataset centered around the open-source OLMo-7B LLM (Groeneveld et al., 2024). Knowing that a pre-training epoch roughly consists of 450k, training data seen before 100k pre-training steps is considered as member, and training data 400k and 450k steps is non-member. Dolma-Book (Zhang & Wu, 2024) also leverages the publicly available pre-training data from OLMo, and samples non-member books from books from the Project Gutenberg [^3] https://www.gutenberg.org/ dated after January 1st, 2024.
@@
Some new datasets discard the static paradigm of existing LLM evaluation datasets, and instead propose a dynamic, periodically updated benchmark to avoid contamination:
  • WikiMIA-24 Fu et al. (2024a) tackle WikiMIA’s as well as any publicly released dataset issue: cutoff dates will need to be adjusted as time goes on. To address the issue, the authors extend WikiMIA’s cutoff to March 2024.
  • ArxivMIA (Liu et al., 2024) gathers paper abstracts from the Computer Science and Maths sections of Arxiv. Abstracts dated after 2024 are considered non-members of LLMs pre-training sets.
  • PatentMIA (Zhang et al., 2024b) constructs a benchmarks where non-members are 5,000 patents from Google-Patents which are in Chinese language and dated after March 1st, 2024, and members are 5,000 such patents dated from before January 1st, 2023.
@@
Scaling datasets used for contamination detection is critical, and such datasets should follow advancements in scale and complexity of datasets used to evaluate LLMs. Samuel et al. (2024) emphasize that as of late 2024, there remains a wide discrepancy between datasets used to evaluate LLM performance and datasets used to detect contamination, with the latter usually lagging the former by several years.
@@ 4.3 Future Challenges
Ethical and legal frameworks (Chang, 2021) are needed to govern the collection, usage, and management of LLM pre-training data. This includes protocols for data privacy, consent, and use that help prevent the incorporation of contaminated data from unethical sources and the contamination of widely used pre-training data sources (e.g., CommonCrawl). Developing contamination detection techniques that do not compromise individual privacy will be crucial (Lin et al., 2010; Hayes & Ohrimenko, 2018).
@@ 5 Conclusion
As LLMs rapidly evolve and the training data become more extensive, LLMs’ performances are inevitably biased by contamination. Eventually, most (if not all) target-answer-based public evaluation datasets end up in commonly used pre-training data dumps, which are updated every few months. Therefore, to help detect and quantify contamination, in this survey, we investigated, structured, and classified the current landscape of contamination detection in LLMs. Depending on a researcher’s level of access to the data and LLM parameters, our survey can efficiently guide toward a set of relevant tools to detect contamination.